# Gradient-based Sample Selection for Faster Bayesian Optimization

### v2

@@ A.2 Theorem 2:Analysis on Regret Bound:
[…]
@@
[…]
@@
which is a confidence bound that applies to $\mathbf{x}^{*}$ as well. For a discretization $\mathcal{X}_{n}$ of size $(\tau_{n})^{d}$, i.e., each coordinate space of $\mathcal{X}_{n}$ has a total of $\tau_{n}$ discrete points, we have the following bound on the closest point $[\mathbf{x}]_{n}$ to $\mathbf{x}$ in $\mathcal{X}_{n}$ to ensure a dense set of discretizations:
@@
Since for all $\mathbf{x}\in\mathcal{X}$, we have $\lim_{n\to\infty}\|\mathbf{x}-[\mathbf{x}]_{n}\|=0$, suggesting that $[\mathbf{x}^{*}]_{n}$ approaches $\mathbf{x}^{*}$ as $n$ increases to infinity. Plugging in, we have:
@@ 1. Regret Bound in the First ${T_{M}}$ Rounds
During the initial $M$ rounds, the strategy relies on the standard GP-UCB. By the well-known GP-UCB regret bounds [46], there exists a constant $C_{1}=\frac{8}{\log(1+\sigma^{-2})}$, pick $\delta\in(0,1)$, and define $\beta_{n}=2\log\left(n^{2}2\pi^{2}/(3\delta)\right)+2d\log\left(n^{2}brd\sqrt{\log(4da/\delta)}\right)$, we have,
@@ B.2 Subset Samples Distribution Study
While BO algorithms are theoretically designed to balance exploitation and exploration, with limited budget in practice, they can over-exploit current best regions before shifting to exploration [49], leading to suboptimal performance in locating the global optimums. With gradient-based sample selection, the relative density of samples near the optimal and suboptimal regions increases, maintaining a more balanced distribution. This subset encourages subsequent iterations to focus on regions outside the optimal and suboptimal regions, promoting the exploration of other regions of the search space. As a result, the over-exploitation issue is mitigated.
@@ B.3 Experiments on Kmeans++ selection
Oglic [42] and Hayakawa [43] proposed to select a subset in RKHS, then employed them to construct the Nyström low-rank approximation. We included it as an additional baseline, experimental results in Fig.8 demonstrate that K-means++ has no clear advantage in cumulative regret compared to our proposed method in subset selection, and it costs more time than our method.
@@ B.4 Experiments on high-dimensional BO methods
We noticed that all these algorithms struggled with high-dimensional tasks, so we included some high-dimensional baselines: REMBO [22] and HESBO [23]. The results are shown in Fig. 9, which shows that the difference between the three methods is subtle on Levy20, but REMBO and HSEBO do not perform well on higher-dimensional tasks because these two dimensionality reduction methods do not use all the dimensional information.
@@ B.5 Experiments on other surrogates
Although our theoretical analysis focuses on Gaussian Process surrogates, our proposed gradient-based sample selection method is not limited to GPs. The core idea of the gradient-based sample selection method is to use gradient information to measure the importance of each sample to model fitting, which is model-independent. Therefore, this method can also be applied to Bayesian neural networks (BNN) [50] and Deep Kernel(DK) methods [51]. We have extended our experiments to other popular surrogate models such as Deep Kernels and Bayesian Neural Networks. Our additional results demonstrate that the proposed sample selection strategy yields performance improvements across these different surrogate models. The updated experimental results can be found in Fig. 11 and Fig. 11.
Figure 10: Cumulative regret of Deep Kernal UCB algorithms on the Hart6 experiment.
Figure 11: Cumulative regret of BNN UCB algorithms on the Hart6 experiment.

### v3

@@ A.2 Theorem 2:Analysis on Regret Bound:
$\beta_{n}\;=\;\frac{\sigma_{\min}(n)\,\Bigl[2\ln\!\frac{4\pi_{n}}{\delta}\;+\;2d\,\ln\!\Bigl(n^{2}\,b\,r\,d\;\sqrt{\ln\!\tfrac{4da}{\delta}}\Bigr)\Bigr]\;-\;A_{n}}{\sigma_{\min}(n)\;+\;B_{n}}\,.$ (22)
@@
$\begin{split}P(r>c)&=\frac{1}{\sqrt{2\pi}}\int_{c}^{\infty}e^{-r^{2}/2}\,dr\\[6.0pt] &=e^{-c^{2}/2}\,\frac{1}{\sqrt{2\pi}}\int_{c}^{\infty}e^{-\frac{(r-c)^{2}}{2}-c(r-c)}\,dr\\[6.0pt] &\leq e^{-c^{2}/2}\,\frac{1}{\sqrt{2\pi}}\int_{c}^{\infty}e^{-\frac{(r-c)^{2}}{2}}\,dr\\[6.0pt] &=e^{-c^{2}/2}\,P(r>0)\\[6.0pt] &=\tfrac{1}{2}\,e^{-c^{2}/2}\,.\end{split}$ (25)
@@
which is a confidence bound that applies to $\mathbf{x}^{*}$ as well. For a discretization $\mathcal{X}_{n}$ of size $(\tau_{n})^{d}$, i.e., each coordinate space of $\mathcal{X}_{n}$ has a total of $\tau_{n}$ discrete points, we have the following bound on the closest point $\mathbf{x}]_{n}$ to $\mathbf{x}$ in $\mathcal{X}_{n}$ to ensure a dense set of discretizations:
@@
Since for all $\mathbf{x}\in\mathcal{X}$, we have $\lim_{n\to\infty}\|\mathbf{x}-[\mathbf{x}]_{n}\|=0$, suggesting that $\mathbf{x}^{*}]_{n}$ approaches $\mathbf{x}^{*}$ as $n$ increases to infinity. Plugging in, we have:
@@ 1. Regret Bound in the First $T_{M}$} Rounds
During the initial $M$ rounds, the strategy relies on the standard GP-UCB. By the well-known GP-UCB regret bounds [Srinivas et al., 2009], there exists a constant $C_{1}=\frac{8}{\log(1+\sigma^{-2})}$, pick $\delta\in(0,1)$, and define $\beta_{n}=2\log\left(n^{2}2\pi^{2}/(3\delta)\right)+2d\log\left(n^{2}brd\sqrt{\log(4da/\delta)}\right)$, we have,
@@ B.2 Subset Samples Distribution Study
While BO algorithms are theoretically designed to balance exploitation and exploration, with limited budget in practice, they can over-exploit current best regions before shifting to exploration [Wang and Ng, 2020], leading to suboptimal performance in locating the global optimums. With gradient-based sample selection, the relative density of samples near the optimal and suboptimal regions increases, maintaining a more balanced distribution. This subset encourages subsequent iterations to focus on regions outside the optimal and suboptimal regions, promoting the exploration of other regions of the search space. As a result, the over-exploitation issue is mitigated.
@@ B.3 Experiments on Kmeans++ selection
Oglic [Oglic and Gärtner, 2017] and Hayakawa [Hayakawa et al., 2023] proposed to select a subset in RKHS, then employed them to construct the Nyström low-rank approximation. We included it as an additional baseline, experimental results in Fig.8 demonstrate that K-means++ has no clear advantage in cumulative regret compared to our proposed method in subset selection, and it costs more time than our method.
@@ B.4 Experiments on high-dimensional BO methods
We noticed that all these algorithms struggled with high-dimensional tasks, so we included some high-dimensional baselines: REMBO [Wang et al., 2016] and HESBO [Nayebi et al., 2019]. The results are shown in Fig. 9, which shows that the difference between the three methods is subtle on Levy20, but REMBO and HSEBO do not perform well on higher-dimensional tasks because these two dimensionality reduction methods do not use all the dimensional information.
@@ B.5 Experiments on other surrogates
Although our theoretical analysis focuses on Gaussian Process surrogates, our proposed gradient-based sample selection method is not limited to GPs. The core idea of the gradient-based sample selection method is to use gradient information to measure the importance of each sample to model fitting, which is model-independent. Therefore, this method can also be applied to Bayesian neural networks (BNN) [Mullachery et al., 2018] and Deep Kernel(DK) methods [Wilson et al., 2016]. We have extended our experiments to other popular surrogate models such as Deep Kernels and Bayesian Neural Networks. Our additional results demonstrate that the proposed sample selection strategy yields performance improvements across these different surrogate models. The updated experimental results can be found in Fig. 10.
Figure 10: Cumulative regret of Deep Kernal and BNN with UCB algorithms on the Hart6 experiment.
B.6 RMSE Analysis of Subset GP Fits
To validate that the selected subset is indeed “representative and informative”, we compare the predictive quality of three surrogate models in terms of RMSE on a test dataset: (i) the full-data GP used in GP-UCB, (ii) the GP trained on the gradient-based subset (GSSBO), and (iii) the GP trained on a randomly chosen subset of the same size (RSSBO). Starting from the 60th iteration, the subset selection algorithm is used. At each iteration, we construct the corresponding GP surrogate and compute the root mean square error (RMSE) between its posterior mean and the true underlying function values on the test set:
$\text{RMSE}=\sqrt{\frac{1}{T}\sum_{i=1}^{T}\left(\mu(x_{i})-f(x_{i})\right)^{2}}.$ (59)
As shown in the figure11, GSSBO consistently achieves RMSE values that are close to, and in many iterations only marginally worse than, the full-data GP, indicating that the gradient-selected subset preserves most of the essential information for accurate function approximation. In contrast, RSSBO exhibits notably higher error and larger variance over time, reflecting that random subset sampling fails to reliably capture the non-redundant information necessary for stable surrogate quality. The fluctuations in all curves are partially due to observation noise and the dynamic subset update process, but GSSBO maintains a relatively low and stable RMSE throughout, which empirically supports our design of combining gradient-based importance with direction-space diversity to build a compact yet informative subset that effectively approximates the full GP.
Figure 11: Cumulative regret of Deep Kernal and BNN with UCB algorithms on the Hart6 experiment.